OpenStudy (anonymous):

3. The antiderivative of the function f(x)=3x^2 + 1/x^1/2 , passing through the point (1; 6) is

OpenStudy (anonymous):

\[f(x) = 3x ^{2} + \frac{ 1 }{ \sqrt{x} }\]

Parth (parthkohli):

Try using the Power Rule of integration...\[\int x^n dx = \dfrac{x^{n+1}}{n+1}\]

OpenStudy (e.mccormick):

These are all about solving for the C. Once you have the antiderivative, which ParthKohli, has given you a great clue for, plug in the (x,y) you were given to find the C. Then you will have a specific antiderivative, rather than the infinite possibilities of them.

Parth (parthkohli):

Yup, but careful, don't mention all that initially. Try to get the asker step-by-step :-)

OpenStudy (anonymous):

okay.. i get its thank u for helping

OpenStudy (e.mccormick):

@ParthKohli Absolutley. It is a process. Get the first part, then go to the next. I find it sometimes helps to clear up that this is more than just solve the antiderivative and move on. Gives people an understanding of why the (x,y) was there and that they can ignore that until they have an antiderivative.